# Variational Supervised Contrastive Learning

### v2

@@ Abstract
Contrastive learning has proven to be highly efficient and adaptable in shaping representation spaces across diverse modalities by pulling similar samples together and pushing dissimilar ones apart. However, two key limitations persist: (1) Without explicit regulation of the embedding distribution, semantically related instances can inadvertently be pushed apart unless complementary signals guide pair selection, and (2) excessive reliance on large in-batch negatives and tailored augmentations hinders generalization. To address these limitations, we propose Variational Supervised Contrastive Learning (VarCon), which reformulates supervised contrastive learning as variational inference over latent class variables and maximizes a posterior-weighted evidence lower bound (ELBO) that replaces exhaustive pair-wise comparisons for efficient class-aware matching and grants fine-grained control over intra-class dispersion in the embedding space. Trained exclusively on image data, our experiments on CIFAR-10, CIFAR-100, ImageNet-100, and ImageNet-1K show that VarCon (1) achieves state-of-the-art performance for contrastive learning frameworks, reaching 79.36% Top-1 accuracy on ImageNet-1K and 78.29% on CIFAR-100 with a ResNet-50 encoder while converging in just 200 epochs; (2) yields substantially clearer decision boundaries and semantic organization in the embedding space, as evidenced by KNN classification, hierarchical clustering results, and transfer-learning assessments; and (3) demonstrates superior performance in few-shot learning than supervised baseline and superior robustness across various augmentation strategies.
@@ 1 Introduction
Ever since its introduction, contrastive learning has become a central paradigm in representation learning, enabling advances across computer vision, natural language processing (NLP), speech, multimodal understanding, and applications in natural sciences [56, 59, 21, 58, 3, 36, 31]. Foundational models such as SimCLR [9], MoCo [27], BYOL [25], and the fully supervised SupCon [32] have enabled capabilities ranging from zero-shot image classification to state-of-the-art sentence embeddings, demonstrating that the simple “pull-together, push-apart” principle scales effectively across domains. Despite their empirical success, these objectives function as heuristic energy functions with opaque statistical meaning. While recent analyses have made progress [28, 41, 48, 42, 55], we still lack a principled account of how contrastive interactions shape embeddings that capture relational structure among data samples. This gap motivates our work to establish a rigorous foundation for contrastive learning through variational inference.
@@
Generative models provide a complementary perspective: they introduce latent variables and estimate parameters via maximum likelihood, assigning explicit probabilistic semantics to the representation space [46]. In Variational Autoencoders (VAEs) [33] and flow-based architectures [37, 11, 19, 54], Euclidean proximity in the latent space corresponds to regions of high data likelihood, endowing distances with a rigorous statistical interpretation. Although contrastive learning is not inherently likelihood-based, recent theory frames InfoNCE as density-ratio maximization: it optimizes the ratio of joint probability over marginal probabilities for positive pairs relative to negatives, prioritizing genuine pairs. Both paradigms ultimately seek embeddings where “near” implies “probable,” though via different routes—generative models integrate over latent variables to maximize likelihood, while contrastive methods directly sculpt the space through positive-negative comparisons. This conceptual convergence invites our unified framework, where variational inference provides probabilistic grounding for contrastive objectives while preserving their proven geometric inductive biases.
@@ 2 Preliminaries
Contrastive Learning. Early metric-learning objectives, the pairwise contrastive loss [12], triplet margin loss [49], and N-pair loss [51], laid the groundwork for modern contrastive frameworks. In contrastive learning, given an input sample $\bm{x}$ with ground-truth class label $r$, an encoder maps $\bm{x}$ to an embedding $\bm{z}\in\mathbb{R}^{d}$, typically $\ell_{2}$-normalized. The field subsequently converged on noise-contrastive formulations like InfoNCE [41]:
@@
where $z_{i}$ and $z_{j}^{+}$ are positive pairs and $\tau$ is a temperature parameter. InfoNCE drives self-supervised systems like SimCLR [9] and MoCo [10, 27], and scales to cross-modal setups in CLIP [43]. For supervised learning, SupCon [32] extends this by treating same-class samples as positives:
@@
where $P(i)$ contains samples of the same class as $i$, and $A(i)$ includes all samples except $i$. Recent theoretical advances [61, 30] and methods like X-CLR [50] further improve these frameworks by analyzing optimization dynamics and introducing graded similarities. Our approach builds on these foundations by reformulating supervised contrastive learning through variational inference while preserving semantic linkages across classes.
Variational Inference. Variational methods provide principled probabilistic frameworks for learning latent representations by maximizing an evidence lower bound (ELBO). Variational Autoencoders (VAEs) [33, 45] pioneered this approach with amortized encoder-decoder architectures learning continuous latent spaces. Later advances improved flexibility through normalizing flows [34, 44] and learned priors, while Bayesian approaches like Bayes-by-Backprop [6] and Monte-Carlo dropout [20] incorporated parameter uncertainty. Recent work has integrated contrastive learning with variational objectives: Noise-Contrastive Prior [2] separates between posteriors and priors to tighten latent fit; variational information-bottleneck augments SimCLR [53]; and Recognition-Parameterised Models reformulate the ELBO to yield InfoNCE-style objectives [1]. These integrations create latent spaces that are simultaneously generative, predictive, and uncertainty-aware. Our approach builds on these foundations by maximizing a class-conditional data likelihood with label-softened targets, preserving contrastive learning’s geometric properties while incorporating explicit probabilistic semantics.
Soft Labeling. Soft labeling replaces hard one-hot targets with distribution-based supervision, first popularized through label-smoothing in Inception-v3 [52] and knowledge distillation [29]. Recent advances employ soft labels to improve representation learning [18, 17]: for noisy targets, smoothed distributions enhance calibration across facial expressions [38] and enlarge inter-class margins in partial-label settings [24]. This concept extends to semi-supervised learning where ESL [39] treats uncertain pixels as soft pseudo-labels, SoftMatch [8] applies confidence-weighted pseudo-labels, and ProtoCon [40] refines labels through online clustering. For cross-modal learning, SoftCLIP [22] strengthens image-text embeddings by weighting related captions. Our approach employs soft labeling as a confidence-adaptive variational distribution that dynamically adjusts class distributions, more uniform for confident samples and more peaked for challenging ones, to enhance representation learning through targeted supervision.
@@ 3 Method
Throughout this study, we use the following notation: Let $\bm{z}\in\mathbb{R}^{d}$ denote the $\ell_{2}$-normalized embedding produced by the encoder (e.g., ResNet-50) for an input sample $\bm{x}$ with ground-truth class index $r\in\{1,\ldots,C\}$. Each class is represented by a unit-norm reference vector $\bm{w}_{r}\in\mathbb{R}^{d}$, obtained by normalizing the class mean: $\bm{w}_{r}=\bar{\bm{z}}_{r}/\|\bar{\bm{z}}_{r}\|_{2}$ where $\bar{\bm{z}}_{r}=|\mathcal{B}_{r}|^{-1}\sum_{i\in\mathcal{B}_{r}}\bm{z}_{i}$ and $\mathcal{B}_{r}=\{\,i\mid r_{i}=r,\,i\in\mathcal{B}\,\}$ denotes the index set of samples from class $r$ in mini-batch $\mathcal{B}$. We set $\tau_{1}$ to be the fixed temperature that scales the logits to adjust the sharpness of the class distribution and $\tau_{2}(\bm{z})$ the confidence-adaptive temperature that softens each sample’s target distribution to regularize learning and control intra-class dispersion. With $\mathbf{\theta}$ being the set of learnable weights of the encoder, $p_{\mathbf{\theta}}(r\mid\bm{z})$ denotes probability the model assigns to the correct class $r$ for embedding $\bm{z}$, $p_{\mathbf{\theta}}(\bm{z}\mid r)$ the class-conditional likelihood of observing $\bm{z}$, and $p_{\mathbf{\theta}}(\bm{z})=\sum_{r^{\prime}}p_{\mathbf{\theta}}(\bm{z}\mid r^{\prime})p(r^{\prime})$, with $r^{\prime}$ the dummy class index, gives the marginal embedding density used in the variational derivation. We introduce $q_{\phi}(r\mid\bm{z})$ as a confidence-adaptive label distribution parameterized by $\phi$, which plays the role of an adaptive temperature for accentuating the probability assigned to the ground-truth class while allocating the remaining mass uniformly across the other $C-1$ classes.
@@ 3.1 The Variational Bound of Class-Conditional Likelihood
To formulate our variational approach, we need to establish an evidence lower bound (ELBO) for the class-conditional likelihood of embeddings. We begin with Bayes’ rule [23, 5], which allows us to express the class-conditional likelihood in terms of the posterior probability and marginal densities:
@@ 4 Experiments
We evaluate our VarCon loss on four standard benchmarks: CIFAR-10, CIFAR-100 [35], ImageNet-100, and ImageNet [47, 15], using the official test splits. Our experiments demonstrate rapid convergence under various training configurations, including different batch sizes, epochs, and hyperparameter settings. For downstream classification, we freeze the pretrained encoder and train only a linear classification layer, and additionally apply KNN-classifier to the embeddings to investigate the learned semantic structure. To isolate the contribution of our learning objective, we employ the same data augmentation techniques and encoder architectures used in previous contrastive learning models [32]: augmentation strategies including SimAugment, AutoAugment [13], and StackedRandAugment [14] with ResNet-50 [26], ResNet-101, ResNet-200, and ViT-Base [16] architectures. We set the learning rate according to the linear-scaling rule $\text{lr}\propto B/256$ with cosine decay. Our best results are achieved with batch sizes of 512 for CIFAR-10/100 (200 epochs), 1,024 for ImageNet-100 (200 epochs), and 4,096 for ImageNet (350 epochs). Throughout, we use SGD with momentum 0.9 and weight decay $10^{-4}$ for smaller datasets including CIFAR-10, CIFAR-100, and ImageNet-100 and LARS [57] optimizer for ImageNet training to ensure stability.
@@ 4.1 Classification Performance
To evaluate our proposed VarCon, we conducted extensive experiments on multiple benchmark datasets for image classification. Table 1 presents performance comparisons against state-of-the-art self-supervised methods (SimCLR [9], MoCo V2 [27, 10], BYOL [25], SwAV [7], VicReg [4], and Barlow Twins [60]) and supervised approaches (standard Cross-Entropy and SupCon [32]). VarCon consistently outperforms all competing methods across all datasets. Compared to SupCon, VarCon achieves 0.43% higher Top-1 accuracy on CIFAR-10 (95.94% vs. 95.51%), 1.72% higher on CIFAR-100 (78.29% vs. 76.57%), 1.28% higher on ImageNet-100 (86.34% vs. 85.06%), and 0.64% higher on ImageNet (79.36% vs. 78.72%). The advantage is more pronounced when compared to self-supervised methods. On ImageNet-100, VarCon surpasses the best self-supervised method (Barlow Twins at 80.83%) by 5.51%, and on ImageNet, it outperforms all self-supervised methods by at least 4.07% in Top-1 accuracy. VarCon also demonstrates superior performance on ImageNet-ReaL (see B.2), a re-annotation with more accurate multi-label ground truth, achieving 84.12% Top-1 accuracy compared to SupCon’s 83.87%. They validate that by explicitly modeling feature uncertainty through our ELBO-derived loss function, VarCon learns more well-defined, generalizable representations.
@@
Self-supervised SimCLR $91.52_{\pm 0.07}$ $99.78_{\pm 0.01}$ $70.67_{\pm 0.12}$ $92.01_{\pm 0.06}$ $71.54_{\pm 0.10}$ $91.56_{\pm 0.05}$ $70.31_{\pm 0.08}$ $90.37_{\pm 0.05}$
MoCo V2 $92.93_{\pm 0.09}$ $99.79_{\pm 0.02}$ $70.01_{\pm 0.09}$ $91.68_{\pm 0.07}$ $78.98_{\pm 0.11}$ $95.20_{\pm 0.03}$ $71.06_{\pm 0.06}$ $90.40_{\pm 0.03}$
BYOL $92.57_{\pm 0.06}$ $99.71_{\pm 0.03}$ $70.50_{\pm 0.11}$ $91.95_{\pm 0.04}$ $80.18_{\pm 0.07}$ $94.86_{\pm 0.05}$ $74.28_{\pm 0.09}$ $91.56_{\pm 0.04}$
@@
VicReg $92.09_{\pm 0.10}$ $99.73_{\pm 0.01}$ $68.51_{\pm 0.08}$ $90.91_{\pm 0.07}$ $79.26_{\pm 0.09}$ $95.06_{\pm 0.04}$ $73.25_{\pm 0.08}$ $91.06_{\pm 0.03}$
Barlow Twins $92.70_{\pm 0.07}$ $99.80_{\pm 0.03}$ $71.02_{\pm 0.11}$ $91.95_{\pm 0.04}$ $80.83_{\pm 0.06}$ $95.24_{\pm 0.05}$ $73.26_{\pm 0.06}$ $91.10_{\pm 0.05}$
Supervised Cross-Entropy $95.07_{\pm 0.08}$ $99.82_{\pm 0.02}$ $74.01_{\pm 0.09}$ $91.89_{\pm 0.06}$ $83.17_{\pm 0.08}$ $95.78_{\pm 0.03}$ $78.20_{\pm 0.07}$ $93.71_{\pm 0.04}$
SupCon $95.51_{\pm 0.06}$ $99.85_{\pm 0.02}$ $76.57_{\pm 0.10}$ $93.50_{\pm 0.05}$ $85.06_{\pm 0.07}$ $96.84_{\pm 0.04}$ $78.72_{\pm 0.06}$ $94.31_{\pm 0.03}$
@@
Table 2: Few-shot learning performance on ImageNet with varying training data availability. We report Top-1 accuracy (mean $\pm$ standard error) on ImageNet using ResNet-50 for both SupCon and our proposed VarCon method across different per-class sample sizes ($N$). Best highlighted in blue.

### v3

@@ Abstract
Contrastive learning has proven to be highly efficient and adaptable in shaping representation spaces across diverse modalities by pulling similar samples together and pushing dissimilar ones apart. However, two key limitations persist: (1) Without explicit regulation of the embedding distribution, semantically related instances can inadvertently be pushed apart unless complementary signals guide pair selection, and (2) excessive reliance on large in-batch negatives and tailored augmentations hinders generalization. To address these limitations, we propose Variational Supervised Contrastive Learning (VarCon), which reformulates supervised contrastive learning as variational inference over latent class variables and maximizes a posterior-weighted evidence lower bound (ELBO) that replaces exhaustive pair-wise comparisons for efficient class-aware matching and grants fine-grained control over intra-class dispersion in the embedding space. Trained exclusively on image data, our experiments on CIFAR-10, CIFAR-100, ImageNet-100, and ImageNet-1K show that VarCon (1) achieves state-of-the-art performance for contrastive learning frameworks, reaching 79.36% Top-1 accuracy on ImageNet-1K and 78.29% on CIFAR-100 with a ResNet-50 encoder while converging in just 200 epochs; (2) yields substantially clearer decision boundaries and semantic organization in the embedding space, as evidenced by KNN classification, hierarchical clustering results, and transfer-learning assessments; and (3) demonstrates superior performance in few-shot learning than supervised baseline and superior robustness across various augmentation strategies. Our code is available at https://github.com/ziwenwang28/VarContrast.
@@ 1 Introduction
Ever since its introduction, contrastive learning has become a central paradigm in representation learning, enabling advances across computer vision, natural language processing (NLP), speech, multimodal understanding, and applications in natural sciences [wang_2024_jojo, yu2023enzyme, gao_etal_2021_simcse, you2020graph, baevski2020wav2vec, laskin2020curl, jia2021scaling]. Foundational models such as SimCLR [SimCLR], MoCo [MoCo_V1], BYOL [BYOL], and the fully supervised SupCon [SupCon] have enabled capabilities ranging from zero-shot image classification to state-of-the-art sentence embeddings, demonstrating that the simple “pull-together, push-apart” principle scales effectively across domains. Despite their empirical success, these objectives function as heuristic energy functions with opaque statistical meaning. While recent analyses have made progress [henaff2020data, oord_2018, saunshi2019theoretical, purushwalkam2020demystifying, isola_wang_2020], we still lack a principled account of how contrastive interactions shape embeddings that capture relational structure among data samples. This gap motivates our work to establish a rigorous foundation for contrastive learning through variational inference.
@@
Generative models provide a complementary perspective: they introduce latent variables and estimate parameters via maximum likelihood, assigning explicit probabilistic semantics to the representation space [rombach2022high]. In Variational Autoencoders (VAEs) [kingma2013auto] and flow-based architectures [lipman2023flow, cheng_2024, fan_2025, wang2025training], Euclidean proximity in the latent space corresponds to regions of high data likelihood, endowing distances with a rigorous statistical interpretation. Although contrastive learning is not inherently likelihood-based, recent theory frames InfoNCE as density-ratio maximization: it optimizes the ratio of joint probability over marginal probabilities for positive pairs relative to negatives, prioritizing genuine pairs. Both paradigms ultimately seek embeddings where “near” implies “probable,” though via different routes—generative models integrate over latent variables to maximize likelihood, while contrastive methods directly sculpt the space through positive-negative comparisons. This conceptual convergence invites our unified framework, where variational inference provides probabilistic grounding for contrastive objectives while preserving their proven geometric inductive biases.
@@ 2 Preliminaries
Contrastive Learning. Early metric-learning objectives, the pairwise contrastive loss [chopra2005learning], triplet margin loss [schroff2015], and N-pair loss [sohn2016improved], laid the groundwork for modern contrastive frameworks. In contrastive learning, given an input sample $\bm{x}$ with ground-truth class label $r$, an encoder maps $\bm{x}$ to an embedding $\bm{z}\in\mathbb{R}^{d}$, typically $\ell_{2}$-normalized. The field subsequently converged on noise-contrastive formulations like InfoNCE [oord_2018]:
@@
where $z_{i}$ and $z_{j}^{+}$ are positive pairs and $\tau$ is a temperature parameter. InfoNCE drives self-supervised systems like SimCLR [SimCLR] and MoCo [chen2021_mocov3, MoCo_V1], and scales to cross-modal setups in CLIP [radford2021learning]. For supervised learning, SupCon [SupCon] extends this by treating same-class samples as positives:
@@
where $P(i)$ contains samples of the same class as $i$, and $A(i)$ includes all samples except $i$. Recent theoretical advances [zhou2024zero, ji2024regcl] and methods like X-CLR [sobal2025mathbbxsample] further improve these frameworks by analyzing optimization dynamics and introducing graded similarities. Our approach builds on these foundations by reformulating supervised contrastive learning through variational inference while preserving semantic linkages across classes. Additional related work is provided in Appendix E.
Variational Inference. Variational methods provide principled probabilistic frameworks for learning latent representations by maximizing an evidence lower bound (ELBO). Variational Autoencoders (VAEs) [kingma2013auto, rezende2014stochastic] pioneered this approach with amortized encoder-decoder architectures learning continuous latent spaces. Later advances improved flexibility through normalizing flows [kingma2016improved, rezende2015variational] and learned priors, while Bayesian approaches like Bayes-by-Backprop [blundell2015weight] and Monte-Carlo dropout [gal2016dropout] incorporated parameter uncertainty. Recent work has integrated contrastive learning with variational objectives: Noise-Contrastive Prior [aneja2021contrastive] separates between posteriors and priors to tighten latent fit; variational information-bottleneck augments SimCLR [wang2022rethinking]; and Recognition-Parameterised Models reformulate the ELBO to yield InfoNCE-style objectives [aitchison2024infonce]. These integrations create latent spaces that are simultaneously generative, predictive, and uncertainty-aware. Our approach builds on these foundations by maximizing a class-conditional data likelihood with label-softened targets, preserving contrastive learning’s geometric properties while incorporating explicit probabilistic semantics.
Soft Labeling. Soft labeling replaces hard one-hot targets with distribution-based supervision, first popularized through label-smoothing in Inception-v3 [szegedy2016rethinking] and knowledge distillation [hinton2015distilling]. Recent advances employ soft labels to improve representation learning [fan2023learnable, fan_generalized_2022]: for noisy targets, smoothed distributions enhance calibration across facial expressions [lukov2022teaching] and enlarge inter-class margins in partial-label settings [gong2024does]. This concept extends to semi-supervised learning where ESL [ma2023enhanced] treats uncertain pixels as soft pseudo-labels, SoftMatch [chen2023softmatch] applies confidence-weighted pseudo-labels, and ProtoCon [nassar2023protocon] refines labels through online clustering. For cross-modal learning, SoftCLIP [gao2024softclip] strengthens image-text embeddings by weighting related captions. Our approach employs soft labeling as a confidence-adaptive variational distribution that dynamically adjusts class distributions, more uniform for confident samples and more peaked for challenging ones, to enhance representation learning through targeted supervision.
@@ 3 Method
Throughout this study, we use the following notation: Let $\bm{z}\in\mathbb{R}^{d}$ denote the $\ell_{2}$-normalized embedding produced by the encoder (e.g., ResNet-50) for an input sample $\bm{x}$ with ground-truth class index $r\in\{1,\ldots,C\}$. Each class is represented by a unit-norm reference vector $\bm{w}_{r}\in\mathbb{R}^{d}$, obtained by normalizing the class mean: $\bm{w}_{r}=\bar{\bm{z}}_{r}/\|\bar{\bm{z}}_{r}\|_{2}$ where $\bar{\bm{z}}_{r}=|\mathcal{B}_{r}|^{-1}\sum_{i\in\mathcal{B}_{r}}\bm{z}_{i}$ and $\mathcal{B}_{r}=\{\,i\mid r_{i}=r,\,i\in\mathcal{B}\,\}$ denotes the index set of samples from class $r$ in mini-batch $\mathcal{B}$. We set $\tau_{1}$ to be the fixed temperature that scales the logits to adjust the sharpness of the class distribution and $\tau_{2}(\bm{z})$ the confidence-adaptive temperature that softens each sample’s target distribution to regularize learning and control intra-class dispersion. With $\mathbf{\theta}$ being the set of learnable weights of the encoder, $p_{\mathbf{\theta}}(r\mid\bm{z})$ denotes probability the model assigns to the correct class $r$ for embedding $\bm{z}$, $p_{\mathbf{\theta}}(\bm{z}\mid r)$ the class-conditional likelihood of observing $\bm{z}$, and $p_{\mathbf{\theta}}(\bm{z})=\sum_{r^{\prime}}p_{\mathbf{\theta}}(\bm{z}\mid r^{\prime})p(r^{\prime})$, with $r^{\prime}$ the dummy class index, gives the marginal embedding density used in the variational derivation. We introduce $q_{\phi}(r\mid\bm{z})$ as a confidence-adaptive label distribution parameterized by $\phi$, which plays the role of an adaptive temperature for accentuating the probability assigned to the ground-truth class while allocating the remaining mass uniformly across the other C1C-1 classes.
@@ 3.1 The Variational Bound of Class-Conditional Likelihood
To formulate our variational approach, we need to establish an evidence lower bound (ELBO) for the class-conditional likelihood of embeddings. We begin with Bayes’ rule [gelman1995bayesian, bishop2006pattern], which allows us to express the class-conditional likelihood in terms of the posterior probability and marginal densities:
@@ 4 Experiments
We evaluate our VarCon loss on four standard benchmarks: CIFAR-10, CIFAR-100 [krizhevsky2009learning], ImageNet-100, and ImageNet [russakovsky2015imagenet, deng2009imagenet], using the official test splits. Our experiments demonstrate rapid convergence under various training configurations, including different batch sizes, epochs, and hyperparameter settings. For downstream classification, we freeze the pretrained encoder and train only a linear classification layer, and additionally apply KNN-classifier to the embeddings to investigate the learned semantic structure. To isolate the contribution of our learning objective, we employ the same data augmentation techniques and encoder architectures used in previous contrastive learning models [SupCon]: augmentation strategies including SimAugment, AutoAugment [cubuk2018autoaugment], and StackedRandAugment [cubuk2020randaugment] with ResNet-50 [he2016deep], ResNet-101, ResNet-200, and ViT-Base [dosovitskiy2020image] architectures. We set the learning rate according to the linear-scaling rule $\text{lr}\propto B/256$ with cosine decay. Our best results are achieved with batch sizes of 512 for CIFAR-10/100 (200 epochs), 1,024 for ImageNet-100 (200 epochs), and 4,096 for ImageNet (350 epochs). Throughout, we use SGD with momentum 0.9 and weight decay $10^{-4}$ for smaller datasets including CIFAR-10, CIFAR-100, and ImageNet-100 and LARS [you2017large] optimizer for ImageNet training to ensure stability. Full hyperparameter settings and optimization details are available in Appendix B.1. Analyses of the training dynamics, such as loss convergence and memory efficiency, are provided in Appendix D.
@@ 4.1 Classification Performance
To evaluate our proposed VarCon, we conducted extensive experiments on multiple benchmark datasets for image classification. Table 1 presents performance comparisons against state-of-the-art self-supervised methods (SimCLR [SimCLR], MoCo V2 [MoCo_V1, chen2021_mocov3], BYOL [BYOL], SwAV [caron2020unsupervised], VicReg [bardes2021vicreg], and Barlow Twins [zbontar2021barlow]) and supervised approaches (standard Cross-Entropy and SupCon [SupCon]). VarCon consistently outperforms all competing methods across all datasets. Compared to SupCon, VarCon achieves 0.43% higher Top-1 accuracy on CIFAR-10 (95.94% vs. 95.51%), 1.72% higher on CIFAR-100 (78.29% vs. 76.57%), 1.28% higher on ImageNet-100 (86.34% vs. 85.06%), and 0.64% higher on ImageNet (79.36% vs. 78.72%). The advantage is more pronounced when compared to self-supervised methods. On ImageNet-100, VarCon surpasses the best self-supervised method (Barlow Twins at 80.83%) by 5.51%, and on ImageNet, it outperforms all self-supervised methods by at least 4.07% in Top-1 accuracy. VarCon also demonstrates superior performance on ImageNet-ReaL (see B.2), a re-annotation with more accurate multi-label ground truth, achieving 84.12% Top-1 accuracy compared to SupCon’s 83.87%. They validate that by explicitly modeling feature uncertainty through our ELBO-derived loss function, VarCon learns more well-defined, generalizable representations.
@@
Self-supervised
SimCLR $1.52_{\pm 0.07} 99$$9.78_{\pm 0.01}$ $70.67_{\pm 0.12}$ $92.01_{\pm 0.06}$ $71.54_{\pm 0.10}$ $91.56_{\pm 0.05}$ $70.31_{\pm 0.08}$ $90.37_{\pm 0.05}$
MoCo V2 $2.93_{\pm 0.09} 99$$9.79_{\pm 0.02}$ $70.01_{\pm 0.09}$ $91.68_{\pm 0.07}$ $78.98_{\pm 0.11}$ $95.20_{\pm 0.03}$ $71.06_{\pm 0.06}$ $90.40_{\pm 0.03}$
BYOL $2.57_{\pm 0.06} 99$$9.71_{\pm 0.03}$ $70.50_{\pm 0.11}$ $91.95_{\pm 0.04}$ $80.18_{\pm 0.07}$ $94.86_{\pm 0.05}$ $74.28_{\pm 0.09}$ $91.56_{\pm 0.04}$
@@
VicReg $2.09_{\pm 0.10} 99$$9.73_{\pm 0.01}$ $68.51_{\pm 0.08}$ $90.91_{\pm 0.07}$ $79.26_{\pm 0.09}$ $95.06_{\pm 0.04}$ $73.25_{\pm 0.08}$ $91.06_{\pm 0.03}$
Barlow Twins $2.70_{\pm 0.07} 99$$9.80_{\pm 0.03}$ $71.02_{\pm 0.11}$ $91.95_{\pm 0.04}$ $80.83_{\pm 0.06}$ $95.24_{\pm 0.05}$ $73.26_{\pm 0.06}$ $91.10_{\pm 0.05}$
Supervised
Cross-Entropy $5.07_{\pm 0.08} 99$$9.82_{\pm 0.02}$ $74.01_{\pm 0.09}$ $91.89_{\pm 0.06}$ $83.17_{\pm 0.08}$ $95.78_{\pm 0.03}$ $78.20_{\pm 0.07}$ $93.71_{\pm 0.04}$
SupCon $5.51_{\pm 0.06} 99$$9.85_{\pm 0.02}$ $76.57_{\pm 0.10}$ $93.50_{\pm 0.05}$ $85.06_{\pm 0.07}$ $96.84_{\pm 0.04}$ $78.72_{\pm 0.06}$ $94.31_{\pm 0.03}$
@@
Table 2: Few-shot learning performance on ImageNet with varying training data availability. We report Top-1 accuracy (mean $\pm$ standard error) on ImageNet using ResNet-50 for both SupCon and our proposed VarCon method across different per-class sample sizes (NN). Best highlighted in blue.